OpenStudy (anonymous):

At 20.0 °C, the volume of hydrogen occupies 12.05 L at 100.1 kPa. What is the new volume if the pressure increases to 120.0 kPa and the temperature remains constant?

OpenStudy (anonymous):

math or physic?

OpenStudy (anonymous):

physics.

OpenStudy (anonymous):

post in physid

OpenStudy (anonymous):

c

OpenStudy (anonymous):

P1V1=P2V2 12.05*100.1=x*120 x=10.05 L

OpenStudy (anonymous):

Thank you @d91196